Your firm is considering an investment that will cost $920,000 today. The investment will produce cash flows of $450,000 in year 1, $270,000 in years 2, 3 and 4, and $200,000 in year 5. The discount rate that your firm uses for projects of this type is 12%. How much would the NPV change if discount rate increases to 8%?

($45,813)

($95,214)

$53,373

$102,774
Answer #1

At 12%:

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=450,000/1.12+270,000/1.12^2+270,000/1.12^3+270,000/1.12^4+200,000/1.12^5

=1094283.98

NPV=Present value of inflows-Present value of outflows

=1094283.98-920,000

=174283.98(Approx)

At 8%%:

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=450,000/1.08+270,000/1.08^2+270,000/1.08^3+270,000/1.08^4+200,000/1.08^5

=1197057.55

NPV=Present value of inflows-Present value of outflows

=1197057.55-920,000

=277057.55(Approx)

Hence change in NPV=277057.55-174283.98

=$102774(Approx).